Find the linearization of the function f(x,y) = ye^x + 2x^2 y at the point (0,2) and use the result to approximate f (0.1, 1.9) to three decimal places.


1
Expert's answer
2021-10-01T15:08:49-0400

"L(x,y)=f(a,b)+\frac{\partial f(x,y)}{\partial x}|_{a,b}(x-a)+\frac{\partial f(x,y)}{\partial y}|_{a,b}(y-b)"


"f(0,2)=2"

"f_x=ye^x+4xy"

"f_x(0,2)=2"

"f_y=e^x+2x^2"

"f_y(0,2)=1"


"L(x,y)=2+2x+y-2=2x+y"


"f(0.1,1.9)=0.2+1.9=2.100"